### v3

@@ 7.4. Characterization of eventually dendric shifts
The proof is almost exactly the same as in the dendric case (see [8, Theorem 5.13]), replacing [8, Proposition 5.7] by [18, Lemma 9.3]. ∎
By [18, Theorem 7.3], if $X$ is eventually dendric and $w$ is of length at least $N$, $\mathcal{D}_{w}(X)$ is on an alphabet of size $1+p_{X}(N+1)-p_{X}(N)$.
@@
7.5. Decidability of eventual dendricity for minimal substitutive shifts
@@
We let $X(\sigma,\tau)$ denote such a shift. In this section, we show that, under minimality, (eventual) dendricity is decidable for this class of shift spaces. Whenever $\tau$ is the identity, the minimality of $X$ is also decidable [11].
It is classical [15] that we can suppose that $\sigma$ is non-erasing ($\sigma(a)\neq\varepsilon$ for every letter) and $\tau$ is a coding ($|\tau(a)|=1$ for every letter) so we always make such an assumption.
Observe that one usually considers the restricted version of morphic shifts consisting in shifts generated by a morphic word. The shift space generated by an infinite word $x\in\mathcal{A}^{\mathbb{N}}$ is the shift space $X_{x}=\{y\in\mathcal{A}^{\mathbb{Z}}\mid\mathcal{L}(y)\subseteq\mathcal{L}(x)\}$. Whenever $x$ is uniformly recurrent (that is, every finite word $u\in\mathcal{L}(x)$ occurs infinitely many times in $x$ and with bounded gaps), the shift space $X_{x}$ is minimal. An infinite word $x$ is morphic if there exist morphisms $\sigma:\mathcal{B}^{*}\to\mathcal{B}^{*}$ and $\tau:\mathcal{B}^{*}\to\mathcal{A}^{*}$ such that $\sigma$ is prolongable on some letter $b\in\mathcal{B}$ (that is, $\sigma(b)\in b\mathcal{B}^{*}$ and $\lim_{n\to+\infty}|\sigma^{n}(b)|=+\infty$) and $x=\tau(\sigma^{\omega}(b))=\lim_{n\to+\infty}\tau(\sigma^{n}(b^{\omega}))$. It is also decidable whether a morphic word is uniformly recurrent [23]. Though, it is not a requirement for it to generate a minimal shift space, i.e., a non uniformly recurrent word $x$ could generate a minimal shift space. Thus, deciding if $X(\sigma,\tau)$ is minimal is an open problem.
@@
The set $B_{g}$ is non-empty by aperiodicity, so let $a\in B_{g}$. For every $n\in\mathbb{N}$, we set $\sigma^{n}(a)=u_{n}a_{n}v_{n}$, where $a_{n}\in B_{g}$ and $u_{n}\in B_{b}^{*}$. By the pigeonhole principle, we can find $k<\ell$ such that $a_{k}=a_{\ell}=b^{\prime}$. We thus have $\sigma^{\ell-k}(b^{\prime})=ub^{\prime}v$ for some word $u\in B_{b}^{*}$, and $\sigma^{n(\ell-k)}(b^{\prime})$ starts with $\sigma^{(n-1)(\ell-k)}(u)\sigma^{(n-2)(\ell-k)}(u)\cdots\sigma^{\ell-k}(u)u$ for all $n$. As $u\in B_{b}^{*}$, the sequence $(\sigma^{n(\ell-k)}(u))_{n\geq 0}$ is eventually periodic. Hence, as $\sigma$ and $\tau$ are non-erasing, $u=\varepsilon$ otherwise this contradicts the aperiodicity of $X(\sigma,\tau)$. In other words, $\sigma^{\prime}=\sigma^{\ell-k}$ is prolongable on $b^{\prime}$ and, by minimality, the word $\tau(\sigma^{\prime\omega}(b^{\prime}))$ generates $X(\sigma,\tau)$. ∎
@@
The existence of $\sigma_{u}$ and $\tau_{u}$ is [23, Proposition 28] and the fact that $\theta_{u}$, $\sigma_{u}$ and $\tau_{u}$ are computable is explained in [23, Section 4]. The bound on the number of possible pairs $(\sigma_{u},\tau_{u})$ is [23, Theorem 29]. ∎
Lemma 58 (Durand [20]).
@@
By Lemma 56, we can compute morphisms $\sigma^{\prime},\tau^{\prime}$ such that $X(\sigma,\tau)=X_{x}$, with $x=\tau^{\prime}(\sigma^{\prime\omega}(b))$. It is decidable whether $x$ is periodic and, if so, one can compute the minimal period [22]. In that case, $X(\sigma,\tau)$ is eventually dendric and the dendricity threshold is bounded by the minimal period. If $x$ not periodic, then we use Theorem 60 and Proposition 61 to conclude. ∎
8. Interval exchanges
An interval exchange transformation is a pair ($(\{I_{a}\mid a\in\mathcal{A}\},T)$, where $\mathcal{A}$ is an alphabet, $\{I_{a}\mid a\in\mathcal{A}\}$ is a partition of $[0,1)$ into left-closed and right-open intervals and $T:[0,1)\to[0,1)$ is a bijection whose restriction to each interval $I_{a}$ is a translation. Another way of defining an interval exchange transformation is by means of a pair $\binom{\leq_{1}}{\leq_{2}}$ of total orders on $\mathcal{A}$. An interval exchange transformation is said to be regular if the orbits of the non-zero left endpoints of the intervals $I_{a}$ are infinite and disjoint. We refer to [27] for more precise definitions.
8.1. Extension graphs in codings of IET
@@
Theorem 63 (Ferenczi-Zamboni [27]).
@@
If $w$ is planar for $(\leq^{L},\leq^{R})$, item 2 is then equivalent to the fact that $\mathcal{E}_{X}(w)$ is a tree. In particular, this proves that codings of regular interval exchanges are dendric shift spaces [6].
@@
8.2. S-adic characterization of regular interval exchange transformations
@@
As in the dendric case, we can reduce the number of vertices of this graph using permutation. As a first step, we can completely fix the right (or top) order. However we can reduce this graph even more. For example, on the alphabet $\{0,1,2,3\}$, the irreducible pairs $\binom{0<1<2<3}{1<3<2<0}$ and $\binom{0<1<2<3}{3<1<0<2}$ are equivalent up to permutation even if they have the same top order. Indeed, $\binom{0<1<2<3}{3<1<0<2}$ is equivalent to $\binom{3<2<1<0}{2<0<1<3}$ which becomes $\binom{0<1<2<3}{1<3<2<0}$ with a simple permutation of the letters. In particular, this then gives a graph with $2$ vertices on a ternary alphabet, as done in [29], and with $9$ vertices for an alphabet of size 4.
9. Future work
@@
Over the alphabet $\mathcal{A}=\{0,1,2\}$, a sufficient [29] set of morphisms is given by the permutations and by
@@
For more general alphabets, it is also possible to find regular languages over the alphabet of elementary morphisms, but only for restricted classes of dendric shifts. Such a regular language is for instance trivially described for Arnoux-Rauzy shifts, and can be obtain from the Rauzy diagrams for regular interval exchange transformations. The Rauzy diagrams encode the behavior of interval exchanges under the Rauzy induction, and this induction process transfers to the codings by means of elementary morphisms [32]. For codings of regular IET, derivation with respect to a word $w$ corresponds to induction on a particular interval associated with $w$, and this induction can be performed by a sequence of Rauzy inductions [17]. Hence the return morphisms appearing in codings of $k$-regular interval exchange form a regular language over the alphabet of elementary morphisms and an automaton can be deduced from the Rauzy diagram.
This raises the question of finding a regular language for the general case. More precisely, if $\mathcal{A}$ is fixed, is there a regular language over the alphabet of elementary morphisms over $\mathcal{A}$ that exactly corresponds to dendric return morphisms over $\mathcal{A}$

### v4

@@ 7.4 Characterization of eventually dendric shifts
The proof is almost exactly the same as in the dendric case (see ((Berthé et al., 2015c, Theorem 5.13))), replacing ((Berthé et al., 2015c, Proposition 5.7)) by ((Dolce and Perrin, 2021, Lemma 9.3)). ∎
By ((Dolce and Perrin, 2021, Theorem 7.3)), if $X$ is eventually dendric and $w$ is of length at least $N$, $\mathcal{D}_{w}(X)$ is on an alphabet of size $1+p_{X}(N+1)-p_{X}(N)$.
@@
7.5 Decidability of eventual dendricity for minimal substitutive shifts
@@
We let $X(\sigma,\tau)$ denote such a shift. In this section, we show that, under minimality, (eventual) dendricity is decidable for this class of shift spaces. Whenever $\tau$ is the identity, the minimality of $X$ is also decidable ((Béal et al., 2021)).
It is classical ((Cobham, 1968)) that we can suppose that $\sigma$ is non-erasing ($\sigma(a)\neq\varepsilon$ for every letter) and $\tau$ is a coding ($|\tau(a)|=1$ for every letter) so we always make such an assumption.
Observe that one usually considers the restricted version of morphic shifts consisting in shifts generated by a morphic word. The shift space generated by an infinite word $x\in\mathcal{A}^{\mathbb{N}}$ is the shift space $X_{x}=\{y\in\mathcal{A}^{\mathbb{Z}}\mid\mathcal{L}(y)\subseteq\mathcal{L}(x)\}$. Whenever $x$ is uniformly recurrent (that is, every finite word $u\in\mathcal{L}(x)$ occurs infinitely many times in $x$ and with bounded gaps), the shift space $X_{x}$ is minimal. An infinite word $x$ is morphic if there exist morphisms $\sigma:\mathcal{B}^{*}\to\mathcal{B}^{*}$ and $\tau:\mathcal{B}^{*}\to\mathcal{A}^{*}$ such that $\sigma$ is prolongable on some letter $b\in\mathcal{B}$ (that is, $\sigma(b)\in b\mathcal{B}^{*}$ and $\lim_{n\to+\infty}|\sigma^{n}(b)|=+\infty$) and $x=\tau(\sigma^{\omega}(b))=\lim_{n\to+\infty}\tau(\sigma^{n}(b^{\omega}))$. It is also decidable whether a morphic word is uniformly recurrent ((Durand, 2013b)). Though, it is not a requirement for it to generate a minimal shift space, i.e., a non uniformly recurrent word $x$ could generate a minimal shift space. Thus, deciding if $X(\sigma,\tau)$ is minimal is an open problem.
@@
The set $B_{g}$ is non-empty by aperiodicity, so let $a\in B_{g}$. For every $n\in\mathbb{N}$, we set $\sigma^{n}(a)=u_{n}a_{n}v_{n}$, where $a_{n}\in B_{g}$ and $u_{n}\in B_{b}^{*}$. By the pigeonhole principle, we can find $k<\ell$ such that $a_{k}=a_{\ell}=b^{\prime}$. We thus have $\sigma^{\ell-k}(b^{\prime})=ub^{\prime}v$ for some word $u\in B_{b}^{*}$, and $\sigma^{n(\ell-k)}(b^{\prime})$ starts with
$\sigma^{(n-1)(\ell-k)}(u)\sigma^{(n-2)(\ell-k)}(u)\cdots\sigma^{\ell-k}(u)u$
for all $n$. As $u\in B_{b}^{*}$, the sequence $(\sigma^{n(\ell-k)}(u))_{n\geq 0}$ is eventually periodic. Hence, as $\sigma$ and $\tau$ are non-erasing, $u=\varepsilon$ otherwise this contradicts the aperiodicity of $X(\sigma,\tau)$. In other words, $\sigma^{\prime}=\sigma^{\ell-k}$ is prolongable on $b^{\prime}$ and, by minimality, the word $\tau(\sigma^{\prime\omega}(b^{\prime}))$ generates $X(\sigma,\tau)$. ∎
@@
The existence of $\sigma_{u}$ and $\tau_{u}$ is ((Durand, 2013b, Proposition 28)) and the fact that $\theta_{u}$, $\sigma_{u}$ and $\tau_{u}$ are computable is explained in ((Durand, 2013b, Section 4)). The bound on the number of possible pairs $(\sigma_{u},\tau_{u})$ is ((Durand, 2013b, Theorem 29)). ∎
Lemma 58 (Durand ((Durand, 1998))).
@@
By Lemma 56, we can compute morphisms $\sigma^{\prime},\tau^{\prime}$ such that $X(\sigma,\tau)=X_{x}$, with $x=\tau^{\prime}(\sigma^{\prime\omega}(b))$. It is decidable whether $x$ is periodic and, if so, one can compute the minimal period ((Durand, 2013a)). In that case, $X(\sigma,\tau)$ is eventually dendric and the dendricity threshold is bounded by the minimal period. If $x$ not periodic, then we use Theorem 60 and Proposition 61 to conclude. ∎
8 Interval exchanges
An interval exchange transformation is a pair ($(\{I_{a}\mid a\in\mathcal{A}\},T)$, where $\mathcal{A}$ is an alphabet, $\{I_{a}\mid a\in\mathcal{A}\}$ is a partition of $[0,1)$ into left-closed and right-open intervals and $T:[0,1)\to[0,1)$ is a bijection whose restriction to each interval $I_{a}$ is a translation. Another way of defining an interval exchange transformation is by means of a pair $\binom{\leq_{1}}{\leq_{2}}$ of total orders on $\mathcal{A}$. An interval exchange transformation is said to be regular if the orbits of the non-zero left endpoints of the intervals $I_{a}$ are infinite and disjoint. We refer to ((Ferenczi and Zamboni, 2008)) for more precise definitions.
8.1 Extension graphs in codings of IET
@@
Theorem 63 (Ferenczi and Zamboni ((Ferenczi and Zamboni, 2008))).
@@
If $w$ is planar for $(\leq^{L},\leq^{R})$, item 2 is then equivalent to the fact that $\mathcal{E}_{X}(w)$ is a tree. In particular, this proves that codings of regular interval exchanges are dendric shift spaces ((Berthé et al., 2015b)).
@@
8.2 S-adic characterization of regular interval exchange transformations
@@
As in the dendric case, we can reduce the number of vertices of this graph using permutation. As a first step, we can completely fix the right (or top) order. However we can reduce this graph even more. For example, on the alphabet $\{0,1,2,3\}$, the irreducible pairs $\binom{0<1<2<3}{1<3<2<0}$ and $\binom{0<1<2<3}{3<1<0<2}$ are equivalent up to permutation even if they have the same top order. Indeed, $\binom{0<1<2<3}{3<1<0<2}$ is equivalent to $\binom{3<2<1<0}{2<0<1<3}$ which becomes $\binom{0<1<2<3}{1<3<2<0}$ with a simple permutation of the letters. In particular, this then gives a graph with $2$ vertices on a ternary alphabet, as done in ((Gheeraert et al., 2022)), and with $9$ vertices for an alphabet of size 4.
9 Future work
@@
Over the alphabet $\mathcal{A}=\{0,1,2\}$, a sufficient ((Gheeraert et al., 2022)) set of morphisms is given by the permutations and by
@@
For more general alphabets, it is also possible to find regular languages over the alphabet of elementary morphisms, but only for restricted classes of dendric shifts. Such a regular language is for instance trivially described for Arnoux-Rauzy shifts, and can be obtain from the Rauzy diagrams for regular interval exchange transformations. The Rauzy diagrams encode the behavior of interval exchanges under the Rauzy induction, and this induction process transfers to the codings by means of elementary morphisms ((Rauzy, 1979)). For codings of regular IET, derivation with respect to a word $w$ corresponds to induction on a particular interval associated with $w$, and this induction can be performed by a sequence of Rauzy inductions ((Dolce and Perrin, 2017)). Hence the return morphisms appearing in codings of $k$-regular interval exchange form a regular language over the alphabet of elementary morphisms and an automaton can be deduced from the Rauzy diagram.
This raises the question of finding a regular language for the general case. More precisely, if $\mathcal{A}$ is fixed, is there a regular language over the alphabet of elementary morphisms over $\mathcal{A}$ that exactly corresponds to dendric return morphisms over $\mathcal{A}$?